# Log-concavity of solutions of parabolic equations related to the Ornstein-Uhlenbeck operator and applications

### v1

@@ 1. Introduction
where $\delta(x)$ is the Dirac distribution.
@@ 3.2. Trotter product formula
By Hille-Yosida theorem, we deduce that the operator $L_{\Omega}$ generates a symmetric and ultracontractive $\mathcal{C}_{0}$-semigroup, denoted by $e^{-tL_{\Omega}}$ (cf. [24]). Using the same arguments as in [11, Chapter 2], we conclude that the semigroup $e^{-tL_{\Omega}}$ can be written in the form
@@
By Trotter product formula [27, formula (3), page vii], we have
@@ 4. Log-concavity of solution of the Fokker-Planck equation
By the spectral mapping theorem [24, Theorem 2.4], we know that the eigenvalues of the self-adjoint operator $L_{\Omega}$ are the eigenvalues of the semigroup $e^{-tL_{\Omega}}$ (which is also a self-adjoint operator) in the Hilbert space $L^{2}(\Omega,\gamma)$, and the corresponding eigenfunctions coincide. So, we have
@@
is uniformly convergent on $\Omega\times\Omega\times[a,+\infty)$ for any $a>0$. Hence (24) holds. Using Proposition 4.4 and [16, Theorem 6.2], it can be proved that the series generated by the first and second partial derivatives are still uniformly convergent. Hence
@@ 6. Appendix
In appendix, we give more details about semigroup theory of Ornstein-Uhlebeck in the class of bounded Lipschitz domains. When we prove the spectral representation of the integral kernel, we assume a priori that the semigroup $e^{-tL_{\Omega}}$ generated by $L_{\Omega}$ has a kernel representation. As previously mentioned in this paper, the reference [26, pages 424-428] has already proved that the Ornstein-Uhlenbeck operator generates an analytic $\mathcal{C}_{0}$-semigroup in the class of bounded Lipschitz domain. In the class of convex bodies, those properties of the semigroup were proved by Cappa [5], and more details were presented in his PhD thesis [6]. We note that Cappa’s proof does not require convexity. The main proof follows the book by Davies [11] and the handbook by W. Arendt [1, Chapter 7], so, we can deduce that the semigroup has a representation as an integral kernel (see for instance [26, Page 422, Theorem 2.18]). Precisely we can retrieve the properties listed below.

### v2

@@ 1. Introduction
where $\delta(x)$ is the Dirac distribution. Very recently, [24] proves the partial log-concavity of kernel when $\Omega$ is bounded and convex, that is, fixed $z\in\Omega$ and $t>0$, $p^{\Omega}_{\gamma}(x,z;t)$ is log-concave in $x\in\Omega$, and the proof relies on an elliptic method. In the present work, we provide a complete answer through an alternative method.
@@ 3.2. Trotter product formula
By Hille-Yosida theorem, we deduce that the operator $L_{\Omega}$ generates a symmetric and ultracontractive $\mathcal{C}_{0}$-semigroup, denoted by $e^{-tL_{\Omega}}$ (cf. [22]). Using the same arguments as in [10, Chapter 2], we conclude that the semigroup $e^{-tL_{\Omega}}$ can be written in the form
@@
By Trotter product formula [26, formula (3), page vii], we have
@@ 4. Log-concavity of solution of the Fokker-Planck equation
By the spectral mapping theorem [22, Theorem 2.4], we know that the eigenvalues of the self-adjoint operator $L_{\Omega}$ are the eigenvalues of the semigroup $e^{-tL_{\Omega}}$ (which is also a self-adjoint operator) in the Hilbert space $L^{2}(\Omega,\gamma)$, and the corresponding eigenfunctions coincide. So, we have
@@
is uniformly convergent on $\Omega\times\Omega\times[a,+\infty)$ for any $a>0$. Hence (24) holds. Using Proposition 4.4 and [15, Theorem 6.2], it can be proved that the series generated by the first and second partial derivatives are still uniformly convergent. Hence
@@ 6. Appendix
In appendix, we give more details about semigroup theory of Ornstein-Uhlebeck in the class of bounded Lipschitz domains. When we prove the spectral representation of the integral kernel, we assume a priori that the semigroup $e^{-tL_{\Omega}}$ generated by $L_{\Omega}$ has a kernel representation. As previously mentioned in this paper, the reference [25, pages 424-428] has already proved that the Ornstein-Uhlenbeck operator generates an analytic $\mathcal{C}_{0}$-semigroup in the class of bounded Lipschitz domain. In the class of convex bodies, those properties of the semigroup were proved by Cappa [5], and more details were presented in his PhD thesis [6]. We note that Cappa’s proof does not require convexity. The main proof follows the book by Davies [10] and the handbook by W. Arendt [1, Chapter 7], so, we can deduce that the semigroup has a representation as an integral kernel (see for instance [25, Page 422, Theorem 2.18]). Precisely we can retrieve the properties listed below.